# A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare

### v1

@@ 3.1 Truthfulness
Ensuring the truthfulness of LLMs in healthcare is vital, as inaccurate information can significantly impact patient care and clinical outcomes. Given their influence on diagnoses and treatment decisions, it is essential to develop effective methods to detect and mitigate hallucinations and factual inaccuracies.
Hallucinations in medical LLMs arise from reliance on unverified sources, biases in training data, and limitations in contextual understanding and sequential reasoning Ahmad et al. (2023). Addressing these issues requires robust evaluation frameworks, self-correction mechanisms, and uncertainty quantification techniques.
The Med-HALT benchmark Pal et al. (2023) is designed to evaluate hallucinations in medical LLMs by using reasoning-based tests like ‘False Confidence’ and ‘None of the Above,’ as well as memory-based tests to assess how well the model recalls medical knowledge. On the other hand, the interactive self-reflection methodology Ji et al. (2023) aims to reduce hallucinations in medical question-answering tasks by introducing an iterative feedback loop where the model refines its responses through self-evaluation and knowledge adjustment.
In the context of multimodal models, Med-HVL Yan et al. (2024b) introduces two key metrics—Object Hallucination and Domain Knowledge Hallucination—to quantify hallucinations in Large Vision-Language Models (LVLMs). This framework also uses the CHAIR (Caption Hallucination Assessment with Image Relevance) metric to assess object hallucinations in image captioning. Med-HallMark Chen et al. (2024) takes it a step further by providing a multi-task evaluation framework with a hierarchical categorization of hallucinations, introducing the MediHall Score for assessing hallucination severity and the MediHallDetector, a multitask-trained LVLM for hallucination detection.
Researchers have also investigated semantic entropy, a probabilistic measure of uncertainty, to detect hallucinations in LLMs. For example, Farquhar et al. (2024) leverage semantic entropy to identify confabulations—hallucinations where the model generates arbitrary or incorrect outputs. While effective, this approach is computationally expensive, limiting its scalability. To overcome this, Han et al. (2024a) introduce Semantic Entropy Probes (SEPs), which approximate semantic entropy directly from hidden states. By eliminating the need for multiple output samples, SEPs significantly reduce computational overhead. Both methods have been successfully applied to biomedical datasets, such as BioASQ.
Although these techniques offer valuable contributions, hallucination mitigation methods often lack adaptability, being either task-specific or requiring expensive retraining. To address this gap, MEDAL Li et al. (2024) introduces a model-agnostic post-processing framework that integrates with any medical summarization model. MEDAL uses a self-examining correction model to improve factual accuracy without adding extra computational costs, providing a practical solution to the issue of medical hallucinations.
Collectively, these studies underscore the multifaceted challenge of ensuring truthfulness in medical LLMs. By leveraging benchmarking frameworks, self-correction mechanisms, and entropy-based uncertainty measures, researchers can develop complementary strategies for detecting, mitigating, and quantifying hallucinations. A key focus of these efforts is the development of quantifiable scoring methods, enabling systematic assessment and comparison across different models. These evaluation techniques not only help identify the most reliable and effective LLMs for healthcare applications but also provide actionable insights for further improvements.
Factual accuracy is fundamental to building trust in LLMs, especially in healthcare, where reliable and verifiable information is critical. However, current LLMs lack effective mechanisms to trace claims back to their original sources, underscoring the urgent need for improved validation techniques to ensure safe and trustworthy medical applications. To address these challenges, several studies have introduced innovative approaches to enhance the transparency, accuracy, and reliability of healthcare LLMs.
@@
Tan et al. (2024) propose an approach that integrates multiple perspectives from scientific literature to evaluate conflicting arguments, thereby improving LLM reasoning. Similarly, Akhtar et al. (2022) introduce PubHealthTab, a table-based dataset designed for validating public health claims against noisy evidence, while Sarrouti et al. (2021) present HEALTHVER, a dataset tailored for evidence-based fact-checking of health-related claims. These structured benchmarks provide a foundation for assessing and refining the reliability of LLM-generated medical information.
Beyond dataset-driven validation, self-correction mechanisms have been explored to improve LLM truthfulness. Gou et al. (2024) introduce CRITIC, a framework inspired by human fact-checking practices, enabling LLMs to validate and refine their responses through iterative feedback and evaluation. Expanding on automated fact-checking, Cohen et al. (2023) propose a cross-examination framework, where an examiner LLM identifies inconsistencies through multi-turn interactions with the original model. Unlike fully automated verification pipelines, CRITIC incorporates human-like evaluation strategies, enhancing the trustworthiness of fact-checking in medical contexts.
Overall, these studies advance the factual accuracy and transparency of LLMs in healthcare by introducing structured benchmarks, iterative validation processes, and automated fact-checking strategies. By incorporating these approaches, researchers can enhance the reliability of medical LLMs, ensuring they deliver more accurate, evidence-based insights to support clinical decision-making.
@@ 3.2 Privacy
LLM-based healthcare applications pose significant privacy risks due to their ability to memorize and reproduce sensitive patient data Das et al. (2024). Unauthorized data exposure can lead to confidentiality breaches, ethical concerns, and compliance violations Pan et al. (2020). Addressing these risks requires privacy safeguards at different stages of model development.
A major challenge is unintended data retention and leakage, where LLMs memorize fine-tuning data, increasing re-identification risks. Studies show that domain-specific LLMs, such as Medalpaca, can retain sensitive data, making privacy breaches more likely Yang et al. (2024a). Additionally, adversarial attacks like prompt injection and inference attacks can further exploit these vulnerabilities, as demonstrated by the SecureSQL benchmark Song et al. (2024).
To mitigate these risks, pre-training privacy safeguards focus on de-identification. Altalla’ et al. (2025) assess GPT-3.5 and GPT-4 in clinical note de-identification and synthetic data generation. Similarly, Liu et al. (2023b) propose a GPT-4-enabled framework for masking private information while maintaining text structure. However, de-identification remains imperfect, as attackers may infer sensitive details from anonymized text.
During fine-tuning, techniques such as federated learning Zhao et al. (2024) and differential privacy Singh et al. (2024), as highlighted by Liu et al. (2024), play a crucial role in safeguarding patient data. Federated learning enables decentralized training without sharing raw data, but it demands high computational resources. Differential privacy adds noise to protect sensitive information but can reduce model accuracy.
Adversarial defenses remain limited. The SecureSQL benchmark Song et al. (2024) highlights LLM vulnerabilities to structured query attacks. While chain-of-thought (COT) prompting offers partial mitigation, it does not eliminate the risk of data exposure.
Researchers address privacy concerns in healthcare LLMs through two primary approaches: de-identification, which alters real data to prevent re-identification, and synthetic data generation, which creates artificial data to eliminate reliance on sensitive patient information. While these strategies enhance privacy protection and maintain model effectiveness, challenges remain in long-term memorization control and adversarial robustness, requiring further research to strengthen data security and prevent unintended information retention.
@@ 3.3 Safety
Ensuring the safety of LLMs in healthcare is critical, as these models must not generate harmful responses. A key safety concern, noted by Han et al. (2024d), is that modifying just 1.1 % of a model’s weights can embed persistent biomedical inaccuracies while maintaining overall performance. This highlights the need for rigorous validation mechanisms and safety assessments to prevent misleading medical information before clinical use.
To systematically assess safety, MedSafetyBench Han et al. (2024b) was introduced as the first benchmark designed to evaluate LLM safety in medical contexts. It includes 1,800 harmful medical queries alongside safety-optimized responses generated using advanced LLMs and adversarial techniques. Results indicate that publicly available medical LLMs fail to meet safety standards, but fine-tuning with MedSafetyBench significantly improves safety without compromising performance.
A major challenge is that adversarial actors can manipulate LLMs to generate unsafe outputs, while excessive safety alignment may induce hallucinations. To address this, Das and Srihari (2024) propose UNIWIZ, a two-step framework that unifies safety alignment and factual knowledge retrieval. Their safety-priming approach synthesizes safety-focused training data, while a retrieval mechanism ensures that model outputs remain factually accurate. Models fine-tuned on UNIWIZ outperform larger state-of-the-art instruction-tuned models across multiple safety and accuracy metrics.
Another key contribution to safety alignment is from Han et al. (2024c), who provide the first comprehensive safety evaluation for medical LLMs (MedLLMs). They define key concepts of medical safety and alignment and introduce Med-Harm, a dataset designed to evaluate both general and medical-specific risks. This dataset assesses how well LLMs handle harmful medical questions, ensuring they adhere to safety and ethical standards in medical AI.
Further advancing safety assessments, Kanithi et al. (2024) introduce MEDIC, a multi-dimensional trustworthiness evaluation framework that systematically assesses medical LLMs across critical dimensions of clinical competence including medical reasoning and clinical safety.
These studies collectively offer a multi-faceted approach to LLM safety in healthcare. MedSafetyBench Han et al. (2024b) provides a standardized benchmark for safety evaluation and fine-tuning, while UNIWIZ Das and Srihari (2024) introduces a structured framework that prevents hallucinations while reinforcing safety. Han et al. (2024c) focus on comprehensive safety alignment, establishing Med-Harm to evaluate domain-specific risks. Lastly, MEDIC Kanithi et al. (2024) offers a holistic evaluation franmework for improving practical application of LLMs in clinical settings. Together, these efforts contribute to a more rigorous and systematic framework for assessing and improving LLM safety in medical applications.
@@ 3.4 Robustness
Enhancing the robustness of LLMs is crucial for their reliability in healthcare applications. A key approach involves developing adversarial test samples tailored to the medical domain, such as synthetic anomaly cases Yuan et al. (2023) and boundary stress testing Wang et al. (2024), to assess model resilience. However, creating clinically meaningful adversarial samples presents unique challenges, as Alberts et al. (2023) highlight the need to align adversarial testing methods with the complexities of real-world medical data, where medical dependencies must be accounted for.
@@
In addition to adversarial testing, uncertainty quantification is another important avenue to improve robustness. LLM-TTA O’Brien et al. (2024) explores test-time adaptation techniques to enhance model performance on rare or unfamiliar cases, which are common in medical diagnostics. Unlike adversarial robustness, which focuses on resistance to manipulated inputs, uncertainty quantification aims to identify when models are likely to be incorrect, providing a complementary safety mechanism.
Another critical question is whether benchmark performance truly reflects medical robustness. MedFuzz Ness et al. (2024) challenges assumptions in MedQA by modifying questions to test if models rely on rigid, dataset-specific patterns rather than genuine clinical reasoning. This research exposes vulnerabilities in LLMs, revealing that subtle changes in input can significantly impact performance, raising concerns about their reliability in dynamic medical environments.
@@
Instruction robustness is also a growing concern. Ceballos-Arroyo et al. (2024) examine how variations in medical instructions affect performance across different LLMs, finding that specialized medical models may be more fragile than general-purpose models when instructions are reworded. This counterintuitive result suggests that excessive domain adaptation may decrease flexibility and reduce robustness.
Adversarial vulnerabilities also pose direct security risks. Yang et al. (2024b) investigate two adversarial attack strategies across medical tasks using real patient data, demonstrating that fine-tuned models are especially vulnerable to poisoning attacks that subtly alter learned weights. While adversarial data does not always degrade general performance, it can introduce dangerous biases into specific medical predictions, making early attack detection a priority.
@@
To protect against adversarial manipulations, Tang et al. (2024) introduce Secure Your Model, a framework that strengthens LLM robustness with cryptographic prompt authentication. This mechanism ensures that only verified and secure prompts are processed, mitigating vulnerabilities associated with prompt injections and adversarial attacks, and reducing the risk of model exploitation in medical contexts.
While all these studies address LLM robustness, they differ in their primary focus. MedFuzz Ness et al. (2024) and Ceballos-Arroyo et al. (2024) expose vulnerabilities in existing benchmarks and task instructions, questioning whether current evaluation methods truly measure robustness or merely reflect dataset biases. In contrast, Yang et al. (2024b) and Alberts et al. (2023) highlight adversarial threats, demonstrating how medical LLMs can be subtly manipulated, raising concerns about their security in real-world applications. Meanwhile, LLM-TTA O’Brien et al. (2024) takes a different approach, focusing on uncertainty quantification rather than adversarial resistance to enhance reliability in handling unfamiliar cases. Secure Your Model Tang et al. (2024) provides an additional security layer by introducing proactive adversarial defenses through prompt protection mechanisms, ensuring resilience against manipulation risks.
These studies highlight that robustness is a multifaceted challenge, requiring advancements in evaluation methods and defensive mechanisms. Ensuring LLMs can handle adversarial scenarios, integrate domain knowledge, and adapt to language variations is crucial for their safe deployment in healthcare. Strengthening robustness through testing and resilience-building enhances the trustworthiness of medical LLMs, making them more reliable in complex clinical settings.
@@ 3.5 Fairness and Bias
Ensuring fairness in LLMs is crucial in healthcare, where biased models can result in unequal treatment outcomes. Research has highlighted biases in clinical data and practice related to race, gender, and disability. For example, Omiye et al. (2023) examine the potential for harmful or inaccurate race-based content in LLMs, while Zack et al. (2024) discusse how language models encode societal biases that can affect healthcare outcomes. These studies underscore the need for fairness in LLM development to ensure equitable healthcare delivery.
Efforts to address bias focus on both detection and mitigation. Swaminathan et al. (2024) offers an automated method for detecting race-based medicine stereotypes, while BiasMedQA Schmidgall et al. (2024) benchmarks cognitive biases in medical tasks across multiple models, revealing varying bias resilience. Mitigation strategies, such as bias education, and one-shot and few-shot bias demonstrations, are proposed to reduce but not fully eliminate bias. Pfohl et al. (2024) introduce frameworks for assessing health equity-related harms in LLMs, including EquityMedQA, a dataset for equity-focused testing. Additionally, Wei et al. (2024) distinguishe between intrinsic fairness, rooted in model training, and behavioral fairness, which relates to model operation in real-world applications, advocating for both to ensure equitable outcomes.
Bias is also explored in closed-source models. Zack et al. (2024) evaluate racial and gender biases in clinical scenarios, finding that models often amplify societal biases. Similarly, Adam et al. (2022) show that biased AI recommendations can affect emergency decisions, while Yang et al. (2024c) identify healthcare disparities in model predictions based on patient demographics. For open-source LLMs, techniques such as reinforcement learning with clinician feedback Zack et al. (2024) and data augmentation during pre-training Parray et al. (2023) enhance training data quality to reduce bias. In contrast, for closed-source models, where internal representations are inaccessible, strategies like instruction fine-tuning Singhal et al. (2023) and prompt engineering Schmidgall et al. (2024) are employed to improve fairness in outputs.
Further, Zhang et al. (2020) investigate how LLM embeddings can encode biases, particularly in clinical tasks, and applies adversarial debiasing to mitigate disparities. Lin and Ng (2023) identify cognitive biases in BERT, while Ke et al. (2024) use a multi-agent framework to explore how LLMs can mitigate biases in clinical decision-making. Additionally, Zhu et al. (2023) introduce CI4MRC, a method that addresses name-related bias in Machine Reading Comprehension (MRC) tasks by applying a causal interventional paradigm.
In addressing bias, solutions vary depending on whether the bias is at the individual level (e.g., name-related information) or dataset level (e.g., biased racial distributions). Open-sourced LLMs benefit from direct accessibility, enabling robust interventions to mitigate bias. In contrast, closed-sourced models, due to their inaccessibility, rely on methods such as instruction fine-tuning and external post-processing tools to refine model outputs and reduce bias.

### v2

@@ 3.1 Truthfulness
Findings in Truthfulness Self-reflection and fact-checking reduce hallucinations but do not scale or generalize well, especially for long-form clinical contexts. Improving truthfulness will require hybrid pipelines that combine retrieval, reasoning, multi-agent self-correction, and uncertainty estimation.
Ensuring the truthfulness of LLMs in healthcare is critical, as inaccurate or fabricated information can directly harm clinical decisions. Hallucinations arise from biased data, weak contextual reasoning, and unverifiable sources Ahmad et al. (2023). Current work targets factual reliability via benchmarking, post-hoc correction, uncertainty quantification, and improved evidence synthesis.
Several benchmarks have emerged to quantify and categorize hallucinations. The Med-HALT benchmark Pal et al. (2023) evaluates hallucination types using reasoning-based tests (e.g., “False Confidence”) and memory checks. In multimodal settings, Med-HVL Yan et al. (2024) distinguishes between Object Hallucination and Domain Knowledge Hallucination.
To mitigate hallucinations, post-hoc correction techniques are gaining traction. MEDAL Li et al. (2024) presents a model-agnostic self-correction module that improves summarization outputs without retraining. Similarly, interactive feedback strategies like self-reflection loops Ji et al. (2023) allow LLMs to iteratively refine their responses.
Uncertainty quantification approaches provide complementary detection tools. Farquhar et al. (2024) apply semantic entropy to flag low-confidence responses, while SEPs Han et al. (2024a) offer a lightweight, hidden-state-based approximation suited for clinical use.
Recent efforts also examine the trustworthiness of evidence synthesis pipelines. Zhang et al. (2024) highlight risks when LLMs generate clinical summaries without grounding, emphasizing the need for transparency in literature retrieval and evidence aggregation. Debate-based evaluation, as explored in MAD Smit et al. (2023), introduces multi-agent deliberation to vet factual consistency in medical QA. Finally, SEND Mohammadzadeh et al. (2024) introduces a neuron dropout technique to detoxify hallucination-prone neurons during training, aiming to improve inherent model truthfulness.
Factual accuracy is critical for trust in healthcare LLMs, where clinical safety relies on reliable, verifiable outputs. Yet, current models often produce ungrounded content and lack source traceability. Recent work addresses this through medical claim benchmarks, self-correction, automated fact-checking, multi-turn verification, and multi-perspective reasoning—advancing transparency, factuality, and clinical relevance.
@@
To support systematic validation, Akhtar et al. (2022) introduce PubHealthTab, a table-based dataset for checking public health claims against noisy evidence, while Sarrouti et al. (2021) propose HEALTHVER, a benchmark for evidence-based fact-checking tailored to medical claims. These resources enable structured evaluation of LLM outputs and form the foundation for improving medical claim verification.
Beyond static benchmarks, dynamic self-correction methods have shown promise. Gou et al. (2024) propose CRITIC, a framework inspired by human fact-checking, in which LLMs iteratively assess and revise their own responses. This process mimics expert reasoning and introduces a layer of critical reflection into model outputs. Complementing this, Cohen et al. (2023) present a cross-examination approach, where a second "examiner" model engages in multi-turn dialogue to probe for factual inconsistencies in the original response. While CRITIC emphasizes human-like evaluation, cross-examination leverages interaction between models to simulate external verification.
To further reduce hallucinations and improve factual consistency, Tan et al. (2024) introduce a method that incorporates multiple scientific perspectives when resolving conflicting arguments, strengthening LLMs’ reasoning capabilities through broader contextual understanding.
Evaluations
Truthfulness is assessed with hallucination/factuality benchmarks (e.g., Med-HALT) and feedback-loop strategies; expert annotations on HealthSearchQA, MedQA, and MultiMedQA are common but costly and subjective.
Limitations
Evaluations remain fragmented: narrow task coverage, varying definitions, closed-source dependencies, and limited generalizability across modalities/clinical domains. Many self-correction methods are task-specific and lack robustness.
@@ 3.2 Privacy
Findings in Privacy LLMs pose serious privacy risks from memorizing and regenerating PHI. Differential privacy and federated learning help but often hurt utility; future work needs fine-grained, instance-level risk estimation across training and inference.
LLMs in healthcare face end-to-end privacy risks due to their tendency to memorize and potentially regenerate sensitive data such as protected health information (PHI) Das et al. (2024); Pan et al. (2020). Key threats include data memorization, insufficient de-identification, and the privacy-utility trade-offs of fine-tuning methods. This section examines current vulnerabilities, mitigation strategies, and emerging approaches for achieving privacy-preserving healthcare LLMs.
Data memorization is a core concern, especially in domain-specific models like Medalpaca Han et al. (2025), which are more likely to retain PHI and pose heightened re-identification risks Yang et al. (2024a). Structured attacks like those demonstrated in SecureSQL Song et al. (2024) reveal that even chain-of-thought (CoT) prompting provides only marginal defense against leakage.
Pre-training privacy measures include de-identification techniques like GPT-4 masking Liu et al. (2023b) and synthetic note generation Altalla’ et al. (2025), though these offer limited protection. Xin et al. (2024) caution that such methods may create a false sense of security, as subtle semantic cues can still lead to PHI leakage.
Fine-tuning methods such as federated learning Zhao et al. (2024a) and differential privacy (DP) Singh et al. (2024) provide stronger safeguards by decentralizing data or adding noise to protect individual records. However, these methods often compromise model performance or scalability Liu et al. (2024a).
Emerging techniques seek to reduce this trade-off. APNEAP Wu et al. (2024) introduces activation patching for privacy neuron editing, reducing leakage without harming utility. Complementarily, Chen and Esmaeilzadeh (2024) offer a broader survey of privacy risks and solutions across generative AI use cases in healthcare.
Ethical and personalization challenges further complicate privacy design. Zhui et al. (2024) emphasize building privacy-conscious frameworks in medical education, while Kirk et al. (2024) caution that overly personalized alignment strategies may inadvertently violate user privacy, advocating instead for bounded personalization.
Evaluations
Typical assessments use membership/attribute inference and reconstruction attacks, plus privacy–utility analyses (e.g., redaction or DP) under controlled settings. Real-world audits are scarce, and proposed risk-prediction or memorization-severity frameworks lack medical-specific benchmarks.
Limitations
Current defenses (e.g., DP, redaction) still trade performance for privacy. Many evaluations assume idealized adversaries, and systematic ways to balance memorization risk and utility—especially for multimodal, long-context models—remain limited.
@@ 3.3 Safety
Findings in Safety Medical LLMs can still produce harmful or misleading content after safety tuning. Benchmarks expose vulnerabilities to adversarial prompts and embedded misinformation. Robust safety demands proactive alignment and multi-stage, clinically grounded evaluations that simulate realistic misuse.
Ensuring safety is critical: small weight edits ($\sim 1.1\%$) can implant lasting biomedical falsehoods without hurting average performance Han et al. (2024d), and poisoning as little as 0.001% of training data can embed persistent misinformation Han et al. (2024e). Key concerns include the ease of injecting persistent falsehoods into model weights, inadequate performance on harmful prompts, trade-offs between safety alignment and hallucination, and privacy-related vulnerabilities that can escalate safety risks. This section explores current benchmarks, safety alignment strategies, and the overlap between safety and privacy threats.
To systematically evaluate harmful outputs, benchmarks like MedSafetyBench Han et al. (2024b) and Med-Harm Han et al. (2024c) use adversarial and real-world queries to test model responses. Results show that even medically fine-tuned LLMs often fail safety criteria unless specifically optimized. MEDIC Kanithi et al. (2024b) broadens this evaluation across dimensions such as reasoning and reliability, offering a holistic safety diagnostic tool.
Safety alignment remains challenging due to its tension with other objectives. UNIWIZ couples safety-driven training with retrieval to reduce unsafe outputs while preserving accuracy Das and Srihari (2024). However, over-alignment increases hallucination, whereas under-alignment permits unsafe behavior, demonstrating the delicate balance required for clinical reliability.
Finally, privacy threats intersect with safety risks. Leemann et al. (2024) show that membership inference attacks, like Gradient Likelihood Ratio (GLiR), can detect whether individual patient data was used in training. This not only violates privacy but also raises safety concerns, as misuse of sensitive information can misguide clinical outcomes.
Evaluations
Methods include manual red teaming, automated stress tests, and healthcare-specific tasks (e.g., drug interactions, diagnostic advice) with expert review, though many prompts still derive from general domains.
Limitations
Mitigations often target generic harms rather than medical-specific risks (e.g., dangerous dosing). Red teaming rarely involves medical experts, and clinically grounded benchmarks with expert-in-the-loop validation remain limited.
@@ 3.4 Robustness
Findings in Robustness LLMs are fragile under distribution shifts, adversarial prompts, and instruction changes. Existing defenses (adversarial testing, test-time adaptation) are often task-specific. Robustness demands context-aware evaluation, multi-agent training, and resilience to real-world perturbations.
Ensuring robustness is vital for clinical deployment. Challenges include adversarial vulnerability, sensitivity to domain shifts and instruction variations, and prompt-based attacks. To address these issues, recent work explores adversarial testing, test-time adaptation, prompt security, data augmentation, and instruction robustness strategies.
Adversarial robustness is addressed through synthetic data generation. Yuan et al. (2023) and Wang et al. (2024) introduce adversarial test samples tailored to the medical domain, such as synthetic anomaly cases and boundary stress testing, to assess model resilience. Alberts et al. (2023) emphasize the importance of aligning adversarial testing methods with real-world medical complexities. In parallel, Gallifant et al. (2024) reveal that simply substituting generic and brand drug names within biomedical benchmarks leads to performance drops of up to 10%, highlighting the fragility of LLMs to clinically trivial lexical shifts.
@@
Uncertainty quantification offers another avenue for robustness. LLM-TTA O’Brien et al. (2024) explores test-time adaptation techniques to enhance model performance on rare or unfamiliar cases, common in medical diagnostics. This approach complements adversarial robustness by identifying instances where models are likely to err.
@@
Instruction robustness is examined by Ceballos-Arroyo et al. (2024), who find that specialized medical models may be more fragile than general-purpose models when instructions are reworded, suggesting that excessive domain adaptation may reduce flexibility.
@@
Prompt security is enhanced by Tang et al. (2024), who introduce a framework that strengthens LLM robustness with cryptographic prompt authentication, mitigating vulnerabilities associated with prompt injections and adversarial attacks.
Data augmentation techniques are employed in MEDSAGE Binici et al. (2025), which uses LLM-generated synthetic dialogues to simulate ASR errors, improving the robustness of medical dialogue summarization systems. Similarly, RobustQA Han et al. (2023) benchmarks the robustness of domain adaptation for open-domain question answering across diverse domains, facilitating the evaluation of ODQA’s domain robustness.
Lastly, prompt engineering strategies, such as out-of-context prompting, are explored by Cotta and Maddison (2024), who demonstrate that applying random counterfactual transformations can improve the fairness and robustness of LLM predictions without additional data or fine-tuning.
Evaluations
Robustness is tested via distribution shifts, adversarial or out-of-domain inputs, synthetic perturbations, and black-box stress tests. Newer setups (e.g., MedQA-Adversarial, RAG robustness tests) probe noisy retrieval and unfamiliar conditions, but clinical realism and standardization remain limited.
Limitations
Lack of standardized, clinically grounded stress tests; overreliance on synthetic or narrow tasks; and brittle/costly mitigations (e.g., adversarial training, RAG) impede deployment. Multimodal and long-context robustness remain underexplored.
@@ 3.5 Fairness and Bias
Findings in Fairness Bias across race, gender, and identity persists in medical LLMs. New benchmarks and mitigations help but are often narrow or misaligned with clinical reality. Progress requires intersectional audits, inclusive datasets, and collaboration with affected communities.
Ensuring fairness is essential: biased outputs can exacerbate disparities in access, diagnosis, and treatment. Key areas of concern include demographic bias (e.g., race, gender, identity), automated detection of these biases, mitigation strategies based on model accessibility, and the need for ethical clarity and conceptual frameworks. Recent work spans benchmark creation, debiasing techniques, prompt interventions, and calls for more transparent fairness evaluations.
Bias identification remains a foundational step. Studies show that LLMs can replicate and even amplify racial, gender, and identity-based biases. For example, Omiye et al. (2023), Zack et al. (2024), and Kim et al. (2023) highlight persistent demographic biases in medical responses. Zhao et al. (2024b) find that diagnostic recommendations vary unfairly by demographic group, while Xie et al. (2024) reveal systematic inequities in outputs concerning race and LGBTQ+ identities. Patel et al. (2024) further demonstrate that LLMs can reinforce social and gender-based stereotypes in sensitive areas such as sexual and reproductive health, underscoring the risks in context-dependent medical interactions.
Detection and benchmarking tools help quantify and monitor these disparities. Swaminathan et al. (2024) propose tools for identifying race-based stereotypes in medical Q&A. Benchmarks such as BiasMedQA Schmidgall et al. (2024), EquityMedQA Pfohl et al. (2024), and FairMedFM Jin et al. (2024) offer frameworks for testing model behavior across diverse patient profiles and clinical contexts.
Mitigation strategies differ by model accessibility. For open-source models, techniques like adversarial debiasing Zhang et al. (2020), causal intervention (CI4MRC) Zhu et al. (2023), multi-agent collaboration Ke et al. (2024), and attribute-neutral modeling Hu et al. (2024) are applied to reduce bias. Data augmentation Parray et al. (2023) and bias-aware embedding assessments Lin and Ng (2023) provide further tools to enhance fairness in pretraining and inference.
Closed-source models present unique challenges due to limited transparency. In these cases, fairness is addressed via instruction fine-tuning Singhal et al. (2023), external prompt engineering Schmidgall et al. (2024), or bounded personalization strategies Kirk et al. (2024), though these are less interpretable and harder to audit.
Ethical and conceptual considerations also play a role. Wei et al. (2024) call for distinguishing between intrinsic and behavioral fairness, while Zhui et al. (2024) and Cotta and Maddison (2024) promote fairness through education and prompt design. Finally, Adam et al. (2022) and Yang et al. (2024b) warn that unchecked bias can distort care decisions and patient trust, emphasizing the stakes of fairness in real-world applications.
Evaluations
Assess disparities across subgroups using quantitative metrics (performance gaps, stereotyping scores) and qualitative audits; new benchmarks (FairMedFM, EquityMedQA) target equity in clinical recommendations, though intersectional analyses remain limited.
Limitations
Many evaluations overlook intersectional/institutional biases, adapt generic NLP methods without clinical causal context, and lack longitudinal assessment. Integration with other trust dimensions (robustness, privacy) is limited.